# Breaking the Curse of Dimensionality: On the Stability of Modern Vector Retrieval

### v1

@@ Appendix B Stability of Filtered Vector Search
Dividing both the numerator and the denominator by $\mathbb{E}[\delta]^{2}$ then gives us
@@
$\geq\frac{\frac{\alpha^{2}p_{c}(1-p_{c})}{\delta_{\max}^{2}}+2\alpha\frac{\mathrm{Cov}(\delta,1_{c})}{\Delta\mathbb{E}[\delta]}}{(1+\frac{\alpha p_{c}}{\mathbb{E}[\delta]})^{2}}$
@@
$\mathrm{RelVar}^{\prime}\geq\frac{\frac{\alpha^{2}p_{c}(1-p_{c})}{\delta_{\max}^{2}}+2\alpha\frac{\mathrm{Cov}(d,1_{c})}{\Delta\mathbb{E}[d]}}{(1+\frac{\alpha p_{c}}{\mathbb{E}[\delta]})^{2}}\geq\frac{\frac{\alpha^{2}p_{c}(1-p_{c})}{\Delta^{2}}-\frac{2\alpha p_{c}}{\Delta}}{(1+\frac{\alpha p_{c}}{\mathbb{E}[\delta]})^{2}}$
@@
$\frac{\alpha^{2}p_{c}(1-p_{c})}{\Delta^{2}}>\frac{2\alpha p_{c}}{\Delta}$
which simplifies to $\alpha>\frac{2\Delta}{1-p_{c}}\geq\frac{2\Delta}{1-p_{\max}}$
@@
Appendix C Stability of Sparse Vector Search
C.1 Proof of Stability

### v2

@@ Appendix D Stability of Filtered Vector Search
Dividing both the numerator and the denominator by $\mathbb{E}[\delta]^{2}$ gives us
@@
$\geq\frac{\frac{\alpha^{2}p_{c}(1-p_{c})}{\Delta^{2}}+2\alpha\frac{\mathrm{Cov}(\delta,1_{c})}{\mathbb{E}[\delta]^{2}}}{(1+\frac{\alpha p_{c}}{\mathbb{E}[\delta]})^{2}}$
@@
$\mathrm{RelVar}^{\prime}\geq\frac{\frac{\alpha^{2}p_{c}(1-p_{c})}{\Delta^{2}}+2\alpha\frac{\mathrm{Cov}(d,1_{c})}{\Delta\mathbb{E}[d]}}{(1+\frac{\alpha p_{c}}{\mathbb{E}[\delta]})^{2}}\geq\frac{\frac{\alpha^{2}p_{c}(1-p_{c})}{\Delta^{2}}-\frac{2\alpha p_{c}}{\mathbb{E}[\delta]}}{(1+\frac{\alpha p_{c}}{\mathbb{E}[\delta]})^{2}}$
@@
$\frac{\alpha^{2}p_{c}(1-p_{c})}{\Delta^{2}}>\frac{2\alpha p_{c}}{\mathbb{E}[\delta]}\geq\frac{2\alpha p_{c}}{\Delta}$
@@
Appendix E Stability of Sparse Vector Search
E.1 Proofs
@@
Remark E.2.
We note out that we state the theorem using $\liminf\mathrm{RelVar}_{m}$ instead of $\lim\mathrm{RelVar}_{m}$. This is convenient because bounding $\liminf$ away from zero is sufficient to prove stability without needing to prove that the relative variance terms converge, which may or may not happen.
E.2 Empirical Validation of Sparse Theorem Assumptions
We empirically validate the assumptions of Theorem 6.4 on $30,522$-dimensional SPLADE-V3 (Lassance et al., 2024) embeddings across several benchmark datasets. We verify concentration of importance, overlap of importance, and finally check the gap condition $Y>X$ required by the theorem.
Concentration of importance.
Overlap of importance.
For random query-document pairs we compute the overlap statistic
$S=\sum_{i\in T_{q}\cap T_{d}}\min\{q_{i}^{p},d_{i}^{p}\}$
Corpus $\alpha$ $\kappa$ $R$ $\rho$ $\gamma$ $\hat{\pi}$ $d_{\max}/d_{\min}$ YXY-X
HotpotQA 0.85 24 3 0.9167 0.00157 0.5194 3.6966 0.0291
TREC-COVID 0.85 24 6 0.9443 0.00236 0.8213 4.7285 2.5830
Natural-Questions 0.85 18 6 0.9537 0.00137 0.5335 2.9501 0.1324
MS MARCO 0.95 18 8 0.9691 0.00049 0.3267 3.2981 0.1159
NFCorpus 0.85 18 8 0.9615 0.0012 0.5412 2.3999 0.1713